# Nearly Zero-Cost Protection Against Mimicry by Personalized Diffusion Models

### v1

@@ Appendix D Discussions
Arbitrary Resolution. As shown in Table 4, FastProtect can effectively handle arbitrary resolution images simply by resizing the MoP to match the resolution of an input image. In contrast, the baseline method, PhotoGuard [salman2023raising], suffers a much greater performance drop compared to ours. There are two main factors that can impact protection performance for arbitrary resolution scenarios. The first factor is the need to adjust the size of the perturbation to match the size of the image when injecting the perturbation. In theory, iterative optimization methods do not face this issue. However, in practice, due to the VRAM constraint of GPUs, images beyond (typically $1024^{2}$-px in a A100) a certain size must be split and protected in segments. On the other hand, our model does not have memory issues, but since the pre-trained MoP is fixed as $512^{2}$-px, it requires resizing to fit an input image. The second factor is the necessity to downsize the protected image due to resolution requirements that each diffusion models has (e.g., $512^{2}$ for SD v1, $768^{2}$ for SD v2, or $1024^{2}$ for SD-XL).
@@
Still, one might wonder about the performance when the protection strength is fixed for all methods. To answer this possible inquiry, we conduct a comparison with all protection frameworks by fixing the protection strength $\eta$ at 8, as shown in Table 7. In most cases, FastProtect demonstrates outstanding results in terms of invisibility while achieving moderate protection efficacy. This balance suggests that FastProtect performs exceptionally well when considering the protection-invisibility trade-off. In the Cartoon domain, Impasto [ahn2024imperceptible] shows better results in invisibility, but our method significantly outperforms Impasto in efficacy, which indicates a favorable trade-off trend.
@@
AdvDM [liang2023adversarial] 0.153 0.267 0.572 357.6 0.438 2.889
PhotoGuard [salman2023raising] 0.107 0.189 0.579 350.9 0.613 3.228
Mist [liang2023mist] 0.129 0.219 0.579 357.0 0.615 3.218
SDST [xue2023toward] 0.152 0.258 0.583 354.1 0.637 3.288
Anti-DB [van2023anti] 0.114 0.184 0.579 347.7 0.452 2.860
Impasto [ahn2024imperceptible] 0.123 0.221 0.579 352.4 0.564 3.141
@@
AdvDM [liang2023adversarial] 0.271 0.440 0.445 212.5 0.579 2.247
PhotoGuard [salman2023raising] 0.209 0.348 0.587 219.1 0.683 2.618
Mist [liang2023mist] 0.223 0.365 0.585 223.7 0.681 2.645
SDST [xue2023toward] 0.237 0.398 0.589 222.7 0.712 2.755
Anti-DB [van2023anti] 0.294 0.468 0.445 225.4 0.429 2.166
Impasto [ahn2024imperceptible] 0.207 0.376 0.602 215.5 0.590 2.294
@@
Limitation & Future Directions. As discussed in Section 5, all protection solutions, including our model, still need improvements in terms of invisibility. However, as analyzed by ahn2024imperceptible, to prevent mimicry, perturbations must be applied broadly across the image. This characteristic makes it very challenging to achieve a high level of invisibility beyond a certain point.
@@ Appendix E Additional Results
In Figure 9, we compare the latency versus input image resolution on both CPU and GPU. On the CPU (Apple silicon), FastProtect completes the protection in only 4 seconds, while all other methods require significantly more time; 15 minutes for PhotoGuard [salman2023raising] and 10 hours for Anti-DreamBooth [van2023anti]. We also include the protection efficacy versus invisibility trade-off comparison for both the Object and Cartoon domains in Figure 10. FastProtect demonstrates an improved trade-off curve in both domains.

### v2

@@ Appendix D Discussions
Arbitrary Resolution. As shown in Table 4, FastProtect can effectively handle arbitrary resolution images simply by resizing the MoP to match the resolution of an input image. In contrast, the baseline method, PhotoGuard [46], suffers a much greater performance drop compared to ours. There are two main factors that can impact protection performance for arbitrary resolution scenarios. The first factor is the need to adjust the size of the perturbation to match the size of the image when injecting the perturbation. In theory, iterative optimization methods do not face this issue. However, in practice, due to the VRAM constraint of GPUs, images beyond (typically $1024^{2}$-px in a A100) a certain size must be split and protected in segments. On the other hand, our model does not have memory issues, but since the pre-trained MoP is fixed as $512^{2}$-px, it requires resizing to fit an input image. The second factor is the necessity to downsize the protected image due to resolution requirements that each diffusion models has (e.g., $512^{2}$ for SD v1, $768^{2}$ for SD v2, or $1024^{2}$ for SD-XL).
@@
Still, one might wonder about the performance when the protection strength is fixed for all methods. To answer this possible inquiry, we conduct a comparison with all protection frameworks by fixing the protection strength $\eta$ at 8, as shown in Table 7. In most cases, FastProtect demonstrates outstanding results in terms of invisibility while achieving moderate protection efficacy. This balance suggests that FastProtect performs exceptionally well when considering the protection-invisibility trade-off. In the Cartoon domain, Impasto [1] shows better results in invisibility, but our method significantly outperforms Impasto in efficacy, which indicates a favorable trade-off trend.
@@
AdvDM [28] 0.153 0.267 0.572 357.6 0.438 2.889
PhotoGuard [46] 0.107 0.189 0.579 350.9 0.613 3.228
Mist [27] 0.129 0.219 0.579 357.0 0.615 3.218
SDST [56] 0.152 0.258 0.583 354.1 0.637 3.288
Anti-DB [51] 0.114 0.184 0.579 347.7 0.452 2.860
Impasto [1] 0.123 0.221 0.579 352.4 0.564 3.141
@@
AdvDM [28] 0.271 0.440 0.445 212.5 0.579 2.247
PhotoGuard [46] 0.209 0.348 0.587 219.1 0.683 2.618
Mist [27] 0.223 0.365 0.585 223.7 0.681 2.645
SDST [56] 0.237 0.398 0.589 222.7 0.712 2.755
Anti-DB [51] 0.294 0.468 0.445 225.4 0.429 2.166
Impasto [1] 0.207 0.376 0.602 215.5 0.590 2.294
@@
Limitation & Future Directions. As discussed in Section 5, all protection solutions, including our model, still need improvements in terms of invisibility. However, as analyzed by Ahn et al. [1], to prevent mimicry, perturbations must be applied broadly across the image. This characteristic makes it very challenging to achieve a high level of invisibility beyond a certain point.
@@ Appendix E Additional Results
In Figure 9, we compare the latency versus input image resolution on both CPU and GPU. On the CPU (Apple silicon), FastProtect completes the protection in only 4 seconds, while all other methods require significantly more time; 15 minutes for PhotoGuard [46] and 10 hours for Anti-DreamBooth [51]. We also include the protection efficacy versus invisibility trade-off comparison for both the Object and Cartoon domains in Figure 10. FastProtect demonstrates an improved trade-off curve in both domains.